1. You prepare a mixture to use to create a colorimetry calibration curve for FeSCN2+ by mixing these things in a 25.00 mL volumetric flask and then diluting to the volume of the flask with 0.3 M HNO3. 5.00 mL 0.00200 M NaSCN 10.00 mL 0.30 M Fe(NO3)3 What is the concentration of FeSCN2+ in this mixture? [Pay attention to the values to determine which reactant is being driven to completely react...]

2. You prepare a mixture to use to measure the equilibrium constant of the reaction of Fe3+ and SCN- ions. You mix these things in a 25.00 mL volumetric flask and then dilute to the volume of the flask with 0.3M HNO3. 3.00 mL 0.00400 M KSCN 10.00 mL 0.00400 M Fe(NO3)3

3.

You prepare a mixture to use to measure the equilibrium constant of the reaction of Fe3+ + SCN- --> FeSCN2+

You mix these things in a 25.00 mL volumetric flask and then diluting to the volume of the flask with 0.3 M HNO3
3.00 mL 0.00400 M KSCN
10.00 mL 0.00400 M Fe(NO3)3

You use colorimetry to determine that, at equilibrium, the FeSCN2+ concentration is 2.4 x 10-5 M.

Create an I-C-E table and find the K for this particular reaction under these particular conditions
Answer #1

1. Fe(NO3)3 is present in large excess; hence, NaSCN is the limiting reactant.

The balanced chemical equation is written as

Fe3+ (aq) + SCN- (aq) ----------> FeSCN2+ (aq)

Since SCN- is the limiting reactant,

[SCN-] = [FeSCN2+]

where the square braces denote molar concentrations.

Use the dilution equation.

M1V1 = M2V2

where M1 = molar concentration of stock NaSCN = 0.00200 M; V1 = volume of stock NaSCN taken = 5.00 mL and V2 = final volume of the solution = 25.00 mL.

M2 = concentration of SCN- in the mixture.

From the dilution equation,

M2 = M1V1/V2

= (0.00200 M)*(5.00 mL)/(25.00 mL)

= 4.00*10-4 M.

Concentration of FeSCN2+ in the mixture = concentration of SCN- = 4.00*10-4 M (ans).

2. Determine the initial concentrations of Fe3+ and SCN- in the test tubes.

[Fe3+]ini = M1V1/V2 = (0.00400 M)*(10.00 mL)/(25.00 mL)

= 1.60*10-3 M

[SCN-]ini = M1V1/V2 = (0.00400 M)*(3.00 mL)/(25.00 mL)

= 4.80*0-4 M.

Set up the ICE chart for the reaction as below.

Fe3+ (aq) + SCN- (aq) <===========> FeSCN2+ (aq)

initial                        1.60*10-3    4.80*10-4                                       -

change                       -x                -x                                                +x

equilibrium          (1.60*10-3 – x)(4.80*10-4 – x)                              x

It is given that x = 2.40*10-5 M.

[Fe3+]eq = (1.60*10-3 – 2.40*10-5) M = 1.576*10-3 M.

[SCN-]eq = (4.80*10-4 – 2.40*10-5) M = 4.56*10-4 M.

[FeSCN2+]eq = 2.40*10-5 M.

Write down the expression for the equilibrium constant K as

K = [FeSCN2+]eq/[Fe3+]eq[SCN-]eq

= (2.40*10-5)/(1.576*10-3)(4.56*10-4) (ignore units since K is dimensionless)

= 33.39 ≈ 33.4 (ans, correct to 3 sig. figs).